OpenStudy (jadedry):

Given that the roots of the equation ( ax^2 +bx +c = 0 ) are beta and (n * beta) show that )((n+1)^2 * ac) =n*b^2 I've figured out that: beta + (n * beta) = -b/a n * beta^2 = c/a (beta + (n * beta) / n*beta^2) = b/c) Unfortunately, that's as far as I got, I'm not sure about to continue. Could someone please provide some hints? Thanks in advance!

imqwerty (imqwerty):

\[(n+1)^2\times ac=n \times b^2\]divide both sides by a^2\[(n+1)^2 \times \frac{ c }{ a }=n \times \left( \frac{ b }{ a } \right)^2\] we know that \[\frac{ c }{a }=n*\beta^2\] and \[\frac{ -b }{ a }=n \beta+\beta\]\[\frac{ b }{ a }=-(n \beta+\beta)\]now just put these two values in the equation u get- \[(n+1)^2 \times n \beta^2 =n \times [-(n \beta+\beta)]^2\] \[(n+1)^2 \times n \beta^2 = n \times [-\beta(n+1)]^2\]\[(n+1)^2 \times n \beta^2 =n \times \beta^2(n+1)^2\]LHS=RHS hence proved

OpenStudy (jadedry):

@imqwerty Thank you! That really clears things up.

imqwerty (imqwerty):

:) no prblm